# Fast Radio Bursts and Electromagnetic Transition Radiation on Gravitational Shockwaves

### v1

@@ Abstract
We solve problem (30), (31) for transition radiation when $\bar{F}_{\mu\nu}$ is EM field of compact source (a magnetar). It is convenient to choose a frame in which the magnetar is at rest at the origin of coordinates after crossing the shockwave front. This means that in the chosen frame, before crossing the front, the magnetar was moving with some velocity. As was shown in [25] , when calculating the data on $u=0$ in the linear order, one can equivalently consider the field of the source located at the origin before crossing the front, but redefine the shockwave profile function $f(x)\to f(x)-f(0)-x^{a}\partial_{a}f(0)$. We will imply that this redefinition in (31) is done.
To find the leading effect in the transition radiation of the magnetar on GSW we approximate the field of the magnetar by the field of a magnetic dipole at the center of coordinates. The corresponding current is $\bar{j}_{\mu}(x)=\delta_{\mu}^{i}\varepsilon_{ijk}M_{j}\,\partial_{k}\delta^{(3)}(\vec{x})$, where $M_{i}$ is a magnetic moment (here Latin indices correspond to spatial coordinates $x,y_{1},y_{2}$). Such a current creates a magnetic field with vector-potential $\bar{A}_{\mu}(x)=\delta_{\mu}^{i}\varepsilon_{ijk}M_{j}\partial_{k}\phi$, $\phi=-1/(4\pi r)$, $r^{2}=x^{2}+y_{i}^{2}$, see [26]. The Cauchy data (31) for the transition radiation are determined by components $\bar{F}_{vi}(x)=(\partial_{v}\mathcal{D}_{i}^{b}-\partial_{i}\mathcal{D}_{v}^{b})\partial_{b}\phi(x)$ ($i$ corresponds to $y_{1}=y,y_{2}=z$), where:
$\begin{split}2\mathcal{D}_{v}^{a}&=M_{y}\delta_{z}^{a}-M_{z}\delta_{y}^{a}~{},~{}\\ \mathcal{D}_{y}^{a}&=-M_{x}\delta_{z}^{a}+2M_{z}\delta_{v}^{a}~{},~{}\\ \mathcal{D}_{z}^{a}&=M_{x}\delta_{y}^{a}-2M_{y}\delta_{v}^{a}~{}.\end{split}$ (32)
It is convenient to use coordinates $(U,r,\Omega)$, where $U=t-r$ and $\Omega$ is a solid angle which parametrizes direction of the unit vector $\vec{e}(\Omega)=(x/r,y^{i}/r)$. The hypersurface $U=0$, $t>0$ is the future light cone of the event when the shockwave hits the magnetar. All time-like trajectories end up inside the cone, $U>0$. By following a method of [25] one can show that distant observers measure the following large $r$ asymptotic at $U>0$
$\hat{F}_{\mu\nu}(U,r,\Omega)\simeq\frac{1}{r}a_{\mu\nu}(U,\Omega)+O(r^{-2})~{},$ (33)
Figure 3: shows EM radiation generated by a null cosmic string passing a magnetar. The notations are identical to those used for Fig. 2. Although the angular distributions are different the pulses of the transition radiation for GSW from the strings and from compact objects have similar forms.
Transition radiation on GSW produced by ultrarelativistic objects and Fast Radio Bursts. As a model of a compact ultrarelativistic object we take a massless point-like particle which moves along the $x$-axis and has the profile function: $f(x)=8GE_{p}\log\sqrt{(y+a)^{2}+z^{2}}$, where $E_{p}$ is the energy of the particle and $a$ is the impact parameter between the particle and the magnetar.
Solution to Cauchy problem (30), (31) for this profile allows us to find asymptotic (33) and the angular distribution of the radiation flux $I(U,\Omega)=r^{2}T_{U}^{r}(U,r,\Omega)$, where $T_{U}^{r}$ are components of the stress-energy tensor of the transition radiation in coordinates $U,r,\Omega$. The results are presented on the Fig. 2. One can point out specific peaks of the intensity, one or two, which appear practically at any observation angle. The form of peaks depend on orientation of the magnetar’s magnetic moment with respect to direction of motion of the GSW. The duration of the peaks is determined by the impact parameter $a$, it as about $0,01~{}a$. The height of the peak is proportional to $(GE_{p}M/a^{3})^{2}$. Fig. 2 shows the dimensionless intensity $\hat{I}(U,\Omega)=(a^{3}/(GE_{p}M))^{2}I(U,\Omega)$. These short pulses of the transition radiation can be the novel engine mechanism for the fast radio bursts as is shown on Fig. 1.
In addition to the impulsive nature of the transition radiation our suggestion is consistent with other requirements. A model of the synchrotron radiation generated by a pulse of magnetic field coming from a magnetar has been formulated by Lyubarsky [3]. According to this model the amplitude of the magnetic field of the pulse should have the form:
$B_{\mathrm{pulse}}=bB_{*}\frac{R_{*}}{r}~{},$ (34)
@@
$\delta t\sim 10^{-2}\ldots 1\,\mathrm{ms}~{}~{},$ (36)
in accord with the requirements for FRB. The ratio $GE_{p}/a$ is not small, for example, if the compact object has a mass comparable to that of a neutron star and moves with a Lorentz factor $10$ at a distance $a\sim 10~{}R_{*}$. In this case $b\sim 10^{-2}$, and the ‘bright spot’ has angle syze $\pi/10$.
It is also interesting to consider ultrarelativistic cosmic strings as alternative type of objects which produce GSW. To simplify things one can take a null infinitely thin cosmic string which is directed along the $z$ axis and moves along the $x$-axis. The corresponding profile function in metric (1) now is $f(x)=8G\pi E_{s}|y+a|$, where $E_{s}$ is the energy of the string per unit length and $a$ is the impact parameter between the string and the magnetar. For infinitely thin null cosmic string, when $\chi(u)=\delta(u)$ in (1), the spacetime geometry is locally flat everywhere except singular points on the string worldsheet. Although the geometry is not formally a GSW shacetime computations are the same. The fact that null cosmic strings passing pulsars generate EM pulses has been first demonstrated in [28] by a similar method based on holonomies of the string spacetime. Here we present computations by using Eqs. (30), (31), see Fig. 3. More useful physical estimates can be found in [27, 28]. The important conclusion from Fig. 3 is that ultrarelativistic cosmic strings passing magnetars also generate pulses of transition radiation with duration (36) and may ensure the FRB engine mechanism.
@@
$I\sim\left(\frac{GE_{p}}{a}\right)^{2}10^{48}\,\mathrm{erg\cdot s^{-1}}~{},$ (37)
for $a\sim 10^{7}$cm. Take a low-energy FRB power of about $10^{38}\,\mathrm{erg\cdot s^{-1}}$ [29] and assume 100% conversion of the transition radiation energy to FRB. Then Eq. (37) implies that $GE_{p}/a\gtrsim 10^{-5}$. For the case of null cosmic strings a similar estimate yields the constraint $GE_{s}\gtrsim 10^{-5}$.
Final remarks. The aim of the present work was to suggest a novel FRB engine mechanism based on the action of gravitational shockwaves on strong magnetic fields. We consider our estimates as tentative, whose purpose was to reveal the basic features of the suggested mechanism. To proceed with quantitative analysis a detailed picture of the EM pulse interaction with the matter surrounding the magnetar is required.
Also we have not discussed other possible physical properties of the transition radiation on GSW. Analysis of scalar models [25] suggests that the transition radiation may include additional EM shocks along the front $u=0$ and along the light cone $U=0$. In particular, if one extends (33) to domain $U<0$ there appear contributions to the right hand side of (33) from contact terms ${\cal J}^{\nu}$ to the Cauchy data. These terms are generated by interaction of the GSW with the magnetar itself and come out as contributions which are non-analytical on $U=0$ (if the magnetar is modelled by a point-like magnetic dipole). We leave analysis of these effects for future work.
The transition radiation is not restricted to electromagnetic fields. A null cosmic string which moves near a compact massive object generates gravitational waves, see [30]. It would be interesting to see if there is a gravitational type transition radiation when gravitational shockwaves perturb gravitational fields of massive objects.
Acknowledgements. The authors are very grateful to Prof. Ilfan Bikmaev for bringing their attention to the problem of FRB, also they thank Dr. Irina Pirozhenko for valuable discussions.

### v2

@@ 5 Transition radiation on GSW produced by ultrarelativistic objects and Fast Radio Bursts
[…]
In addition to the impulsive nature of the transition radiation our suggestion is consistent with other requirements. A model of the synchrotron radiation generated by a pulse of magnetic field coming from a magnetar has been formulated by Lyubarsky Lyubarsky:2014jta . According to this model the amplitude of the magnetic field of the pulse should have the form:
$B_{\mathrm{pulse}}=bB_{*}\frac{R_{*}}{r}\,,$ (42)
@@
$\delta t\sim 10^{-2}\ldots 1\,\mathrm{ms}\,,$ (44)
in accord with the requirements for FRB. The ratio $GE_{p}/a$ is not small, for example, if the compact object has a mass comparable to that of a neutron star and moves with a Lorentz factor 101010 at a distance $a\sim 10~R_{*}$. In this case $b\sim 10^{-2}$, and the ‘bright spot’ has angle size $\pi/10$.
It is also interesting to consider ultrarelativistic cosmic strings as alternative type of objects which produce GSW. To simplify things one can take a null infinitely thin cosmic string which is directed along the $z$ axis and moves along the $x$-axis. The corresponding profile function in metric (1) now is
$f(x)=8G\pi E_{s}|y+a|\,,$ (45)
where $E_{s}$ is the energy of the string per unit length and $a$ is the impact parameter between the string and the magnetar. For infinitely thin null cosmic string, when $\chi(u)=\delta(u)$ in (1), the spacetime geometry is locally flat everywhere except singular points on the string worldsheet. Although the geometry is not formally a GSW spacetime computations are the same. The fact that null cosmic strings passing pulsars generate EM pulses has been first demonstrated in Fursaev:2023lxq by a similar method based on holonomies of the string spacetime. Here we present computations by using Eqs. (32), (33), see figure 3. More useful physical estimates can be found in Fursaev:2022ayo ; Fursaev:2023lxq . The important conclusion from figure 3 is that ultrarelativistic cosmic strings passing magnetars also generate pulses of transition radiation with duration (44) and may ensure the FRB engine mechanism.
Figure 3: shows EM radiation generated by a null cosmic string passing a magnetar. The notations are identical to those used for figure 2. Although the angular distributions are different the pulses of the transition radiation for GSW from the strings and from compact objects have similar forms.
@@
$I\sim\left(\frac{GE_{p}}{a}\right)^{2}10^{48}\,\mathrm{erg\cdot s^{-1}}\,,$ (46)
for $a\sim 10^{7}$cm. Take a low-energy FRB power of about $10^{38}\,\mathrm{erg\cdot s^{-1}}$ Bochenek:2020zxn and assume 100% conversion of the transition radiation energy to FRB. Then Eq. (46) implies that $GE_{p}/a\gtrsim 10^{-5}$. For the case of null cosmic strings a similar estimate yields the constraint $GE_{s}\gtrsim 10^{-5}$.
6 Final remarks
The aim of the present work was to suggest a novel FRB engine mechanism based on the action of gravitational shockwaves on strong magnetic fields. We consider our estimates as tentative, whose purpose was to reveal the basic features of the suggested mechanism. To proceed with quantitative analysis a detailed picture of the EM pulse interaction with the matter surrounding the magnetar is required.
Also we have not discussed other possible physical properties of the transition radiation on GSW. It has long been established Lichnerowicz1 ; Lichnerowicz2 ; Lichnerowicz3 ; Modugno:1976uf ; Modugno:1979rn that a plane gravitational shockwave generates an accompanying electromagnetic shockwave at $u=0$. However, in addition to this shock, as well as the outgoing radiation flux described in our work, other physical effects can potentially be observed. The analysis of scalar models Fursaev:2024czx suggests that the transition radiation may include additional EM shock along the light cone $U=0$. In particular, if one extends (39) to domain $U<0$ there appear contributions to the right hand side of (39) from contact terms ${\cal J}^{\nu}$ to the Cauchy data. These terms are generated by interaction of the GSW with the magnetar itself and come out as contributions which are non-analytical on $U=0$ (if the magnetar is modelled by a point-like magnetic dipole). We leave analysis of these effects for future work.
The transition radiation is not restricted to electromagnetic fields. A null cosmic string which moves near a compact massive object generates gravitational waves, see Fursaev:2023oep . It would be interesting to see if there is a gravitational type transition radiation when gravitational shockwaves perturb gravitational fields of massive objects.
Acknowledgements.
The authors are very grateful to Prof. Ilfan Bikmaev for bringing their attention to the problem of FRB, also they thank Dr. Irina Pirozhenko for valuable discussions.